OpenStudy (anonymous):

What is the slope-intercept form of: 5x-9y=-7

OpenStudy (anonymous):

http://www.okc.cc.ok.us/maustin/EquationsofLine/Equations%20of%20a%20Line.htm I think you'll thank me later ! :)

jhonyy9 (jhonyy9):

y=(5x+7)/9

OpenStudy (anonymous):

\[\LARGE 5x-9y=-7\] \[\LARGE -9y=-5x-7 \quad \quad \quad |(-1)\] \[\LARGE 9y=5x+7 \quad \quad \quad |:9\] \[\LARGE \frac{9y}{9}=\frac{5x}{9}+\frac{7}{9} \] \[\LARGE y=\frac{5}{9}x+\frac{7}{9} \]